Going skiing. Leon drove 270 miles to the lodge in the same time as pat drove 330 miles to the lodge. If Pat drove 10 miles per hour faster than Leon, then how fast did each of them drive?
Answer by Paul(988)

EQUATION:
330(x)=270(x+10)
330x-270x=2700
60x=2700
x=45 (Leon's speed)
45+10=55(Pat's speed)
Paul.